OpenStudy (anonymous):

1,1,2,3,5,8,13 next three numbers and rule

OpenStudy (anonymous):

fibonacci...

OpenStudy (anonymous):

n-2+n-1=n

OpenStudy (anonymous):

\[y _{n}=x _{n-1}+x _{n-2}\]

OpenStudy (anonymous):

waylander, your equation doesnt work the same because your using series notation in an equation....

OpenStudy (anonymous):

:P I was trying to type it before you posted yours... apologies yours is the correct answer

OpenStudy (anonymous):

lmao its not a race... though i did speed up when i saw you typing XDDDD

OpenStudy (anonymous):

Oh it's totally a race on the mathematics part here. On the other areas perhaps we can be slow and ponderous and think about our answers. Here in Mathematics we need to be quick and silent killers

OpenStudy (anonymous):

Oh and tell me how to do fractions!

OpenStudy (rosey):

u should b a fan of me :)

OpenStudy (anonymous):

\[F(n) = \frac{1}{\sqrt{5}}[(\frac{1+\sqrt{5}}{2})^{n}-(\frac{1-\sqrt{5}}{2})^{n}]\] This gives you the nth Fibonnaci number. Tell your teacher this, i bet he will be shocked lol.

OpenStudy (anonymous):

im wondering if it would be a good idea to add a diagram of how to solve it using pascals triangle

OpenStudy (anonymous):

A Q-matrix, even....